Show that hydrogen peroxide (H2O2) is a liquid at 25°C by using the data below.

ΔHf° (kJ mol-1) data: H2O2(l) = –187.8; H2O2(g) = –136.3

So (J mol-1 K-1) data: H2O2(l) = 110; H2O2(g) = 233
Answer #1

H2O2 (l) -----------> H2O2 (g)

ΔH°rxn = ΔHf° products - ΔHf° reactants

             = - 136.3 - (- 187.8)

             = 51.5 kJ

ΔSorxn = 233 - 110

            = 123 J/K

ΔGo = ΔH - T ΔS

0 = ΔH - T ΔS

T = ΔH / ΔS = 51500 / 123

T = 418.7 K

boiling point = 145 oC

the estimated boiling point is more than room temperature. so

this is liquid .